# SEA: Low-Resource Safety Alignment for Multimodal Large Language Models via Synthetic Embeddings

### v1

@@ 4.4 Experimental Validation
We conduct experiments to show that VA-Safetybench is a challenging safety benchmark. The experiments are conducted across four video-based MLLMs including Qwen2-VL-7b Wang et al. (2024b), Qwen2-VL-2B, VideoLLaMA2-7B Cheng et al. (2024), and VideoLLaMA2.1-7B, and four audio-based MLLMs including Qwen2-Audio Chu et al. (2024), Qwen-Audio Chu et al. (2023), SALMONN-7B Tang et al. (2024), and SALMONN-13B. Following the evaluation protocol established in Liu et al. (2024c), we use GPT-4o-mini as the evaluator. The evaluation metric is Attack Success Rate (ASR), calculated as the percentage of responses flagged as unsafe within the evaluation set.
Tables 1 and 2 present the experimental results. The findings can be summarized as follows:
@@
VA-SafetyBench poses significant challenges. Both video and audio-based MLLMs demonstrate a high ASR. In video MLLMs, the best case is the Qwen2-VL-7b, which achieved a 71.33% ASR in DM+TYPO. While the ASR for DM and TYPO is generally lower than that of DM+TYPO, we found that in many instances, the MLLMs failed to correctly interpret the content of the videos, leading to safe outputs. Therefore, as the performance of MLLMs improves in the future, DM and TYPO may pose even greater threats.
For audio MLLMs, the highest ASR is found in SALMONN-13B at 64.34%. KPT is generally higher than Audio-only, indicating that distributing harmful instructions across text and audio better activate model’s toxicity. KPT + noise generally performs better than KPT, indicating that using noise for interference or hiding harmful information in the time dimension makes it easier to bypass safety mechanisms.
@@
Table 3: We present experimental results for an image-based MLLM (Llava-v1.5-7b-hf), with separate evaluation on the safety benchmark and benchmarks on general capabilities. The results on safety benchmarks are presented on the left of the vertical line, with lower scores indicating better performance. The results on benchmarks of general capability are presented on the right of the vertical line, with higher scores indicating better performance. Bold values indicate the best performance.
@@
Table 4: The experimental results conducted on Video-based MLLMs (Qwen2-VL-7b).
Approaches VASafetyBench (%) MVBench VideoMME
@@
Table 5: The experimental results conducted on Audio-based MLLMs (Qwen2-Audio-7b).
Approaches VASafetyBench (%) AIRBench
@@ 5.1 Experimental Setup
Backbones. We select the widely used MLLM backbone for each modality: LLava-v1.5-7b-hf Liu et al. (2024b) for images, Qwen2-VL-7b Wang et al. (2024b) for videos, and Qwen2-Audio-7b Chu et al. (2024) for audio.
Baselines. For image-based MLLMs, we have three baselines: (1) VLGuard Zong et al. (2024) utilizes 2k harmful and 1k harmless image-text pairs for SFT alignment. (2) Textual SFT uses 3k textual alignment samples and SFT as the alignment method. (3) Textual DPO utilizes 3k textual samples and DPO as the alignment method. Since there is no related work on safe alignment, only the two baselines, Textual SFT and Textual DPO, are used for video and audio modalities.
Training Datasets. Following the settings in Hu et al. (2024), we sample 3k examples from SafeRLHF Ji et al. (2024a), including 2k harmful samples and 1k harmless samples, as training data for SEA, Textual SFT, and Textual DPO. There is a slight difference in the data format. VLGuard consists of image-text pairs, whereas our training data only contain plain text data.
Evaluation Benchmark. For safety assessment, we employ MM-SafetyBench for image-based MLLMs and VA-SafetyBench for video and audio-based MLLMs. To evaluate general capabilities, we utilize MMMU Hendrycks et al. (2021) and POPE Li et al. (2023) for image models, MVBench Li et al. (2024a) and VideoMME Fu et al. (2024) for video models, and AIR-Bench Yang et al. (2024b) for audio models.
@@
Implementation Details. We conducted embedding optimization training of SEA on a single RTX 3090 GPU. All MLLMs were set to a maximum of 100 training epochs. The learning rates for LLava-v1.5, Qwen2-VL, and Qwen2-Audio were set to 0.02, 0.02, and 0.05, respectively, with cosine annealing updates. For efficiency, we implement an early stopping mechanism during embeddings optimization by evaluating embedding quality every 10 gradient update steps. For style control, optimization is considered successful if the output is the same with the ground truth. For content control, given that the harmful phrase is categorized as “activity”, we consider optimization successful if $N-1$ out of $N$ words match, allowing for minor verb tense variations while preserving semantic meaning. Given that the harmful phrase is categorized as “product”, we require exact word-for-word matching. Training is terminated early when both content and style optimizations are successful. For failed optimization samples, we directly use the embeddings from the last epoch.
For the safety alignment training, we implemented both SFT and DPO training strategies for SEA. More details about the experimental setup can be found in Appendix B.
@@ 5.2 Main results
Comparison between SEA and Textual Alignment. Both Textual SFT and Textual DPO belong to Textual Alignment approaches. SEA demonstrates comparable safety capabilities to Textual Alignment methods under Text-only attacks while significantly lowering the ASR of multimodal attacks. For instance, in the most challenging SD-TYPO task, SEA SFT’s ASR decreased by 34.95% compared to Textual SFT, and SEA DPO’s ASR decreased by 29.64% compared to Textual DPO.
Compared to models without safety alignment, Textual SFT and Textual DPO effectively reduce the ASR of text-only attacks. Still, their effectiveness against multimodal attacks is limited, which is particularly evident in the image-based MLLM (Llava-v1.5) and audio-based MLLM (Qwen2-Audio).
For general performance, whether using SFT or DPO training strategies, the overall performance of SEA is close to that of Textual Alignment. In summary, compared to Textual Alignment, SEA can significantly reduce the safety risks introduced by additional modalities without sacrificing overall performance.
Comparison between embeddings and physical multimodal data. VLGuard is a baseline trained on physical image-text pairs while SEA operates on embeddings and textual instructions. Since both VLGuard and SEA SFT employ SFT as their alignment method and use equivalent training dataset sizes, it is a fair comparison. The overall performance of SEA SFT is comparable to that of VLGuard, but SEA SFT demonstrates significantly higher safety capabilities against multimodal attacks than VLGuard. The underlying reason may be that SEA SFT always optimizes the image embeddings that the model deems most aligned with harmful content, resulting in higher data quality. In contrast, images collected from the real world may not be correctly interpreted by the model regarding harmful content.
Comparison between DPO and SFT. The SFT-based approaches demonstrate stronger security than the DPO-based methods, but they typically results in a general performance decline. This is because using the reference model in DPO aids in maintaining overall performance. Notably, aside from a slight degradation in the Qwen2-VL-7b, the general performance of SEA DPO in other model does not decline compared to the original model. Therefore, we recommend using DPO as the training strategy for SEA.
@@ 5.3 Efficiency and Quality of Embedding Optimization
To validate the efficiency of the embedding optimization, we recorded the optimization success rate (OSR) and the time consumed for embedding optimization across 3k samples. To check whether the model consistently believes that the optimized embeddings contain information from the content control samples, we designed three rewritten versions of the content control instruction, such as “Could you explain what is occurring in the image?”, and used them to query the MLLMs regarding the content in the optimized embeddings. The proportion of model outputs containing the specified content out of total samples is reported as the validation success rate (VSR).
Table 6: The OSR, average time consumption, and VSR of the embedding optimization on three models.
Models OSR(%) Average Time(s) VSR(%)
Llava-v1.5-7b-hf 98.17 23.86 87.76
@@
Table 6 presents the statistical results. SEA successfully finds embeddings for specified content and style in more than 93% of the cases across all models, demonstrating good generalization even when faced with instructions not seen in the content control samples, indicating that the embeddings are of high quality. Furthermore, each sample requires an optimization time of no more than 24 seconds on a single RTX 3090 GPU, which is significantly lower than the cost of manually collecting real data. Since optimization is performed on individual samples, SEA allows for parallel embedding optimization of a large-scale textual dataset across multiple GPUs, further saving computational time.
@@ 6 Conclusions
The high cost of constructing multimodal datasets poses a significant challenge to developing safety alignment. In this paper, we demonstrate that synthetic embeddings can substitute for real additional model data, allowing for effective multimodal safety alignment relying solely on text. The high performance exhibited across multiple MLLMs, including images, videos, and speech, validates the universality of the proposed SEA method. Until high-quality, large-scale real multimodal datasets are released, it holds promise as a safety solution for emerging MLLMs.
@@
Although SEA has shown promising performance, it remains a preliminary exploration. Currently, we have only used two samples to optimize the embeddings, lacking attempts to further enhance embedding quality with more complex designs. Moreover, this work’s exploration of SEA is limited to safety alignment. Simple modifications to SEA’s data preparation process would suffice for other alignment tasks, such as improving honesty. Its effectiveness will need to be explored in future work.
@@ A.1 Motivation
The proposal of VA-SafetyBench. This work explores low-resource safety alignment methods generally applicable to various modal MLLMs, including image, video, and audio MLLMs. Quantifying the safety risks of models is crucial for evaluating alignment performance. However, there is currently a lack of safety benchmarks for video-based MLLMs. Their projects are still in a partially developed stage for the two existing security evaluation works Ying et al. (2024); Yang et al. (2024a) of audio-based MLLMs. Therefore, VA-SafetyBench needs to be constructed to help evaluate SEA’s performance.
@@
Reduction of Scenarios MM-SafetyBench contains 13 scenarios, but we have excluded political lobbying, legal opinion, financial advice, health consultation, and government decision-making, as these safety considerations are somewhat overly stringent. For instance, financial advice tests the ability of MLLMs to refuse to provide financial opinions. Similar data is difficult to find in existing alignment training datasets, which poses challenges for evaluating algorithms.
@@ A.2 Prompts for Scene Descriptions Generation
Models DM+TYPO (Temporal) DM+TYPO (Spatial)
@@ A.3 Comparison of Temporal Stitching and Spatial Stitching in DM+TYPO
In the DM+TYPO task of VA-SafetyBench, the videos generated by the model and the TYPO videos are stitched along the timeline. In MM-SafetyBench, the Stable Diffusion images and TYPO images are connected in pixel space. In fact, similar operations can also be performed in videos, such as adding TYPO subtitles to the bottom of each video frame. We have compared the experimental results of the two stitching methods in Table 7. Across all models, the temporally stitched datasets exhibited higher levels of harmfulness, so we adopted this approach.
A.4 The complete benchmark evaluation results.
The complete evaluation results of VA-SafetyBench expanded by scenarios are presented in Tables 8 and 9.
@@
Scenarios Qwen2-VL-7b Qwen2-VL-2B
@@
Illegal Activitiy 0 36.08 46.39 84.54 41.75 3.09 11.34 30.93 81.44 31.7
Hate Speech 0 8.59 28.22 61.96 24.69 6.75 3.68 33.13 61.96 26.38
Malware Generation 6.82 22.73 56.82 84.09 42.61 22.73 25 45.45 79.55 43.18
Physical Harm 11.81 30.56 53.47 82.64 44.62 32.64 38.89 46.53 84.03 50.52
Economic Harm 13.93 9.84 24.59 33.61 20.49 20.49 10.66 23.77 31.97 21.72
Fraud 0.65 18.18 46.75 82.47 37.01 13.64 13.64 38.31 81.82 36.85
Sex 13.76 4.59 32.11 47.71 24.54 31.19 4.59 28.44 50.46 28.67
Privacy Violence 1.44 30.94 37.41 92.09 40.47 15.11 17.27 28.78 83.45 36.15
Average 6.05 20.18 40.72 71.13 34.52 18.20 15.63 34.41 69.33 34.39
Scenarios VideoLLaMA2-7B VideoLLaMA2.1-7B
@@
Illegal Activitiy 4.12 44.33 25.77 59.79 33.50 1.03 46.39 65.98 72.16 46.39
Hate Speech 10.43 14.72 16.56 35.58 19.32 4.29 12.88 34.36 42.33 23.46
Malware Generation 29.55 27.27 31.82 47.73 34.09 9.09 27.27 54.55 65.91 39.20
Physical Harm 29.86 34.72 27.78 55.56 36.98 11.81 33.33 56.25 58.33 39.93
Economic Harm 10.66 9.84 6.56 22.95 12.50 12.3 11.48 20.49 26.23 17.62
Fraud 14.94 30.52 31.82 55.84 33.28 1.95 29.22 53.9 62.99 37.01
Sex 12.84 2.75 4.59 13.76 8.48 14.68 2.75 16.51 21.1 13.76
Privacy Violence 12.95 34.53 21.58 47.48 29.13 2.88 31.65 47.48 69.06 37.76
Average 15.66 24.83 20.81 42.33 25.90 7.25 24.37 43.69 52.26 31.89
@@
Scenarios Qwen2-Audio-7b-Instruct Qwen-Audio-Chat
Text-only Audio-only KPT KPT+noise Average Text-only DM TYPO DM+TYPO Average
Illegal Activitiy 1.03 2.06 40.21 16.49 14.94 0 37.11 73.2 57.73 42.01
Hate Speech 1.84 1.84 20.86 15.34 9.97 2.45 50.31 43.56 34.36 32.67
Malware Generation 31.82 15.91 54.55 38.64 35.23 34.09 50 75 65.91 56.25
Physical Harm 23.61 1.25 49.31 39.58 28.43 22.92 58.33 68.06 54.17 50.87
Economic Harm 16.39 16.39 23.77 21.31 19.46 14.75 27.87 30.33 29.51 25.61
Fraud 5.19 3.25 44.16 22.73 18.83 3.25 54.55 68.83 53.9 45.13
Sex 20.18 13.76 12.84 14.68 15.36 10.09 42.2 33.94 31.19 29.35
Privacy Violence 17.27 7.19 28.78 23.74 19.24 10.07 57.55 75.54 50.36 48.38
Average 14.66 7.70 34.31 24.06 20.18 12.20 47.24 58.55 47.14 41.28
Scenarios SALMONN-7B SALMONN-13B
Text-only DM TYPO DM+TYPO Average Text-only DM TYPO DM+TYPO Average
Illegal Activitiy 16.49 - 51.55 80.41 49.48 15.46 - 73.2 84.54 57.73
Hate Speech 31.29 - 32.52 62.58 42.13 25.15 - 53.37 64.42 47.64
Malware Generation 75 - 31.82 70.45 59.09 77.27 - 52.27 59.09 62.87
Physical Harm 69.44 - 54.17 86.81 70.14 59.03 - 66.67 77.78 67.82
Economic Harm 22.95 - 22.95 36.07 27.32 28.69 - 28.69 35.25 30.87
Fraud 57.14 - 51.95 76.62 61.90 70.13 - 65.58 80.52 72.07
Sex 27.52 - 15.6 32.11 25.07 36.7 - 24.77 27.52 29.66
Privacy Violence 52.52 - 68.35 82.73 67.86 56.83 - 76.26 85.61 72.9
Average 44.04 - 41.11 65.97 50.37 46.15 - 55.10 64.34 55.19

### v3

@@ 4.4 Experimental Validation
The experiments are carried out on four video-based MLLMs including Qwen2-VL-7b Wang et al. (2024b), Qwen2-VL-2b, VideoLLaMA2-7b Cheng et al. (2024), and VideoLLaMA2.1-7b, and four audio-based MLLMs including Qwen2-Audio Chu et al. (2024), Qwen-Audio Chu et al. (2023), SALMONN-7b Tang et al. (2024), and SALMONN-13b. Following the evaluation protocol established in Liu et al. (2024c), we use GPT-4o-mini as the evaluator. The evaluation metric is Attack Success Rate (ASR), calculated as the percentage of responses flagged as unsafe within the evaluation set.
Tables 1 and 2 present the experimental results. The findings can be summarized as follows.
@@
VA-SafetyBench poses significant challenges. Both video and audio-based MLLMs demonstrate high ASRs. For video-based MLLMs, the best case is the Qwen2-VL-7b, which achieved a 71.33% ASR in DM+TYPO. While the ASR for DM and TYPO respectively is generally lower than that of DM+TYPO, we found that in many instances, the MLLMs failed to correctly interpret the content of the videos, leading to safe outputs. Therefore, as the performance of MLLMs improves in the future, DM and TYPO may pose even greater threats. For audio-based MLLMs, the highest ASR is found in SALMONN-7b at 65.97%. KPT is generally higher than Audio-only, indicating that distributing harmful instructions across text and audio better activate model’s toxicity. KPT + Noise generally performs better than KPT, indicating that using noise for interference or hiding harmful information in the time dimension makes it easier to bypass safety mechanisms.
@@
Table 3: We present experimental results for an image-based MLLM (LLaVA-1.5-7b-hf), with separate evaluation on the safety benchmark and benchmarks of general capabilities. The results on safety benchmarks are presented on the left of the vertical line, with lower scores indicating better performance. The results on benchmarks of general capability are presented on the right of the vertical line, with higher scores indicating better performance. Bold values indicate the best performance.
@@
Table 4: The experimental results conducted on a video-based MLLM (Qwen2-VL-7b).
Approaches VA-SafetyBench (%) MVBench VideoMME
@@
Table 5: The experimental results conducted on an audio-based MLLM (Qwen2-Audio-7b).
Approaches VA-SafetyBench (%) AIRBench
@@ 5.1 Experimental Setup
Backbones. We select the widely used MLLM backbone for each modality: LLaVA-1.5-7b-hf Liu et al. (2024b) for images, Qwen2-VL-7b Wang et al. (2024b) for videos, and Qwen2-Audio-7b Chu et al. (2024) for audio.
Baselines. For image-based MLLMs, we have three baselines: (1) VLGuard Zong et al. (2024) utilizes 2k harmful and 1k harmless image-text pairs for SFT alignment. (2) Textual SFT (DPO) uses 3k textual samples for SFT (DPO) alignment. (3) GM SFT (DPO) uses a text-driven generative model (GM) to synthesize additional modal data for 3k textual samples, guided by the content guiding texts of SEA.
Since there is no related work on safe alignment, only the two baselines, Textual SFT (DPO) and GM SFT (DPO), are used for video and audio modalities. The generative models used for the three modalities are FLUX.1-dev  [^2] https://github.com/black-forest-labs/flux, CogVideoX-2b Yang et al. (2024c), and ChatTTS  [^3] https://github.com/2noise/ChatTTS, which differ from the models used in the benchmark construction.
Training Datasets. Following the settings in Hu et al. (2024), we sample 3k examples from the textual alignment dataset SafeRLHF Ji et al. (2024a), including 2k harmful samples and 1k harmless samples, as training dataset for SEA, Textual SFT (DPO), and GM SFT (DPO). For details on sampling, please refer to Appendix B.2. VLGuard’s training data is based on the dataset provided in the original work Zong et al. (2024), most of which consist of real-world image-text pairs.
Evaluation Benchmark. For safety assessment, we employ MM-SafetyBench for image-based MLLMs and VA-SafetyBench for video and audio-based MLLMs. To evaluate general capabilities, we utilize MMMU Hendrycks et al. (2021) and POPE Li et al. (2023) for image-based MLLMs, MVBench Li et al. (2024a) and VideoMME Fu et al. (2024) for video-based MLLMs, and AIR-Bench Yang et al. (2024b) for audio-based MLLMs.
@@
Implementation Details. We conducted embedding optimization training of SEA on a single RTX 3090 GPU. All MLLMs were set to a maximum of 100 training epochs. The learning rates for LLaVA-1.5, Qwen2-VL, and Qwen2-Audio were set to 0.02, 0.02, and 0.05, respectively, with cosine annealing updates. For efficiency, we implement an early stopping mechanism during embeddings optimization by checking whether the optimization is successful every 10 gradient update steps. For style control, optimization is considered successful if the output includes the complete guiding text. For content control, when the harmful phrase is categorized as “activity”, we consider optimization successful if $N-1$ out of $N$ words match, allowing for minor verb tense variations while preserving semantic meaning. When the harmful phrase is categorized as “product”, we require exact word-for-word matching. Training is terminated early when both content and style optimizations are successful. For failed optimization samples, we directly use the embeddings from the last epoch. For the safety alignment training, we implemented both SFT and DPO training strategies for SEA. More details about the experimental setup can be found in Appendix B.
@@ 5.2 Main Results
Comparison between SEA and textual alignment. Both Textual SFT and Textual DPO belong to textual alignment approaches. Compared to models without safety alignment, Textual SFT and Textual DPO effectively reduce the ASR of text-only attacks. Still, their effectiveness against multimodal attacks is limited, which is particularly evident in the image-based MLLM (LLaVA-1.5) and audio-based MLLM (Qwen2-Audio). SEA demonstrates comparable safety capabilities to textual alignment methods under Text-only attacks while significantly lowering the ASR of multimodal attacks. For instance, in the most challenging SD-TYPO task, SEA SFT’s ASR decreased by 34.95% compared to Textual SFT, and SEA DPO’s ASR decreased by 29.64% compared to Textual DPO. For general performance, whether using SFT or DPO training strategies, the overall performance of SEA is close to that of textual alignment methods. In summary, compared to textual alignment methods, SEA can significantly reduce the safety risks introduced by additional modalities without sacrificing general performance.
Comparison between SEA embeddings and physical multimodal data. In the baselines, VLGuard and GM SFT (DPO) were trained on physical image-text pairs. The results indicate that, except for SEA DPO being slightly inferior to GM DPO in the audio modality, SEA consistently demonstrated better safety performance in every comparison group. Additionally, SEA’s general performance was comparable to these baselines. Further observations of the data generated by the generative model revealed that the audio modality produced the highest quality output, with minimal information loss when converting text to speech, while the generated images and videos often show lower relevance to the guiding text. This may explain GM DPO’s strong performance in the audio modality. On the other hand, the audio modality conveys information through spoken voice, which has a stronger capacity to represent harmful content. Images and videos communicate harmful information through abstract visual concepts, which can lead to MLLMs not correctly interpreting them. This might be one of the reasons why VLGuard and GM SFT (DPO) perform worse than SEA in these two modes. In contrast, SEA consistently captures embeddings that MLLMs perceive as highly relevant to the guiding text, demonstrating stable performance across all modalities. This validates the potential of SEA for future applications in new modal MLLMs.
Comparison between DPO and SFT. The SFT-based approaches demonstrate stronger security than the DPO-based methods, but they typically results in a general performance decline. This is because using the reference model in DPO aids in maintaining general performance. Notably, aside from a slight degradation in the Qwen2-VL-7b, the general performance of SEA DPO in other model does not decline compared to the original model. Therefore, we recommend using DPO as the training strategy for SEA.
@@ 5.3 Efficiency and Quality of Embedding Optimization
To validate the efficiency of the embedding optimization, we recorded the optimization success rate (OSR) and the time consumed for embedding optimization across 3k samples. To check whether the model consistently believes that the optimized embeddings contain information from the content control samples, we designed three rewritten versions of the content control instruction, such as “Could you explain what is occurring in the image?”, and used them to query the MLLMs regarding the content in the optimized embeddings. The proportion of model outputs containing the target content out of total samples is reported as the Generalization Success Rate (GSR).
Table 6 presents the statistical results. SEA successfully finds embeddings for specified content and style in more than 93% of the cases across all models, demonstrating good generalization even when faced with instructions not seen in the content control samples, indicating that the embeddings are of high quality. The cases in Appendix B.4 further validates this. It should be acknowledged that some optimization failures and low-quality embeddings still exist. Since the SEA embeddings are optimized solely based on gradients from the MLLMs, the quality of embedding optimization depends on the model’s internal knowledge. As a result, embedding optimization is prone to failure in domains not covered by the model’s knowledge. However, if the MLLM inherently lacks the relevant knowledge, even an unaligned model would struggle to provide harmful guidance for related malicious instructions. Therefore, retaining low-quality embeddings does not compromise safety, but it may degrade general performance. This might explain the general performance decline of SEA DPO in the video modality. To address this, future improvements could include introducing a filtering mechanism based on embedding quality to discard low-quality embeddings before safety alignment training.
In terms of efficiency, each sample requires an optimization time of no more than 24 seconds on a single RTX 3090 GPU, which is significantly lower than the cost of manually collecting real data. Since optimization is performed on individual samples, SEA allows for parallel embedding optimization of a large-scale textual dataset across multiple GPUs, further saving computational time.
Due to space limitations, additional experiments and analyses regarding SEA can be found in Appendix B.3 and B.4.
Table 6: The OSR, average time consumption, and GSR of the embedding optimization on three models.
Models OSR(%) Average Time(s) GSR(%)
LLaVA-1.5-7b-hf 98.17 23.86 87.76
@@ 6 Conclusions
The high cost of constructing multimodal datasets poses a significant challenge to developing safety alignment. In this paper, we demonstrate that synthetic embeddings can substitute for real additional model data, allowing for effective multimodal safety alignment relying solely on text. The high performance demonstrated by MLLMs across various modalities such as images, videos, and audio, validates the universal applicability of the proposed SEA method. Before the release of high-quality, large-scale real multimodal datasets, it holds promise as a safety solution for emerging MLLMs.
@@
Although SEA has shown promising performance, it still has limitations. On one hand, the optimization of SEA embeddings relies on the model’s own knowledge, making it susceptible to failures on harmful concepts that are not covered by the model. Given the diversity of the world, SEA is likely to fail on concepts which are more unique to different cultures of the world (e.g. a threat in a local dialect or a particular type of food which is poisonous in a country). Since the model is unlikely to provide accurate guidance for concepts it doesn’t understand, filtering out these failed embeddings before safety alignment training seems like a viable solution. However, it remains crucial for the model to correctly perceive these harmful concepts and responsibly reject related instructions. Finding a way to achieve this in low-resource settings is still an open question.
On the other hand, while style control enhances the diversity of SEA embeddings, it still cannot guarantee coverage of a sufficiently broad range of harmful scenarios. Further strategies for enhancing diversity, such as designing more samples for optimization, still need to be explored. We leave these two issues for future works.
@@
Acknowledgments
This work was supported by National Natural Science Foundation of China through grant 62406114, 62306117 and 62472181, the Fundamental Research Funds for the Central Universities through grant 2024ZYGXZR074, Guangdong Basic and Applied Basic Research Foundation through grant 2025A1515011413 and 2024A1515010220, Local Science and Technology Development Fund of Hebei Province Guided by the Central Government of China through grant 246Z0102G, the “Pioneer” and “Leading Goose” R&D Program of Zhejiang through grant 2025C02044, Hebei Natural Science Foundation through grant F2024210008, the Guangzhou Basic and Applied Basic Research Foundation through grant 2024A04J3681, GJYC program of Guangzhou through grant 2024D03J0005, National Key R & D Project from Minister of Science and Technology through grant 2024YFA1211500, and South China University of Technology-TCL Technology Innovation Fund.
@@ A.1 Motivation
The Proposal of VA-SafetyBench. This work explores low-resource safety alignment methods generally applicable to various modal MLLMs, including image, video, and audio MLLMs. Quantifying the safety risks of models is crucial for evaluating alignment performance. However, there is currently a lack of safety benchmarks for video-based MLLMs. For audio-based MLLMs, the projects of two existing works Ying et al. (2024); Yang et al. (2024a) are still not yet fully developed. Therefore, it is necessary to establish VA-SafetyBench to help assess the performance of SEA.
@@
Reduction of Scenarios MM-SafetyBench contains 13 scenarios, but we have excluded political lobbying, legal opinion, financial advice, health consultation, and government decision-making, as these safety considerations are somewhat overly stringent. For instance, the financial advice scenario tests the ability of MLLMs to refuse to provide financial opinions. Similar data is difficult to find in existing alignment training datasets, which poses challenges for evaluating algorithms.
@@ A.2 Prompts for Scene Descriptions Generation
Models DM+TYPO (Spatial) DM+TYPO (Temporal)
@@ A.3 Comparison of Temporal Stitching and Spatial Stitching in DM+TYPO
In the DM+TYPO task of VA-SafetyBench, the videos generated by the model and the TYPO videos are stitched along the timeline. In MM-SafetyBench, the Stable Diffusion images and TYPO images are connected in pixel space. In fact, similar operations can also be performed in videos, such as adding TYPO subtitles to the bottom of each video frame. We have compared the experimental results of the two stitching methods in Table 7. Across all models, the temporally stitched datasets exhibited higher levels of harmfulness, so we adopted this stitching approach.
A.4 The Complete Benchmark Evaluation Results
Tables 8 and 9 present the complete evaluation results of VASafetyBench, including the results for each scenario.
@@ A.4 The Complete Benchmark Evaluation Results
Scenarios Sample size Qwen2-VL-7b Qwen2-VL-2b
@@
Illegal Activitiy 97 0 36.08 46.39 84.54 41.75 3.09 11.34 30.93 81.44 31.70
Hate Speech 163 0 8.59 28.22 61.96 24.69 6.75 3.68 33.13 61.96 26.38
Malware Generation 44 6.82 22.73 56.82 84.09 42.61 22.73 25.00 45.45 79.55 43.18
Physical Harm 144 11.81 30.56 53.47 82.64 44.62 32.64 38.89 46.53 84.03 50.52
Economic Harm 122 13.93 9.84 24.59 33.61 20.49 20.49 10.66 23.77 31.97 21.72
Fraud 154 0.65 18.18 46.75 82.47 37.01 13.64 13.64 38.31 81.82 36.85
Sex 109 13.76 4.59 32.11 47.71 24.54 31.19 4.59 28.44 50.46 28.67
Privacy Violence 139 1.44 30.94 37.41 92.09 40.47 15.11 17.27 28.78 83.45 36.15
Average - 6.05 20.18 40.72 71.13 34.52 18.20 15.63 34.41 69.33 34.39
Scenarios Sample size VideoLLaMA2-7b VideoLLaMA2.1-7b
@@
Illegal Activitiy 97 4.12 44.33 25.77 59.79 33.50 1.03 46.39 65.98 72.16 46.39
Hate Speech 163 10.43 14.72 16.56 35.58 19.32 4.29 12.88 34.36 42.33 23.46
Malware Generation 44 29.55 27.27 31.82 47.73 34.09 9.09 27.27 54.55 65.91 39.20
Physical Harm 144 29.86 34.72 27.78 55.56 36.98 11.81 33.33 56.25 58.33 39.93
Economic Harm 122 10.66 9.84 6.56 22.95 12.50 12.30 11.48 20.49 26.23 17.62
Fraud 154 14.94 30.52 31.82 55.84 33.28 1.95 29.22 53.90 62.99 37.01
Sex 109 12.84 2.75 4.59 13.76 8.48 14.68 2.75 16.51 21.10 13.76
Privacy Violence 139 12.95 34.53 21.58 47.48 29.13 2.88 31.65 47.48 69.06 37.76
Average - 15.66 24.83 20.81 42.33 25.90 7.25 24.37 43.69 52.26 31.89
@@
Scenarios Sample size Qwen2-Audio-7b-Instruct Qwen-Audio-Chat
Text-only Audio-only KPT KPT+noise Average Text-only Audio-only KPT KPT+noise Average
Illegal Activitiy 97 1.03 2.06 16.49 40.21 14.94 0 37.11 57.73 73.20 42.01
Hate Speech 163 1.84 1.84 15.34 20.86 9.97 2.45 50.31 34.36 43.56 32.67
Malware Generation 44 31.82 15.91 38.64 54.55 35.23 34.09 50.00 65.91 75.00 56.25
Physical Harm 144 23.61 1.25 39.58 49.31 28.43 22.92 58.33 54.17 68.06 50.87
Economic Harm 122 16.39 16.39 21.31 23.77 19.46 14.75 27.87 29.51 30.33 25.61
Fraud 154 5.19 3.25 22.73 44.16 18.83 3.25 54.55 53.90 68.83 45.13
Sex 109 20.18 13.76 14.68 12.84 15.36 10.09 42.20 31.19 33.94 29.35
Privacy Violence 139 17.27 7.19 23.74 28.78 19.24 10.07 57.55 50.36 75.54 48.38
Average - 14.66 7.70 24.06 34.31 20.18 12.20 47.24 47.14 58.55 41.28
Scenarios Sample size SALMONN-7b SALMONN-13b
Text-only Audio-only KPT KPT+noise Average Text-only Audio-only KPT KPT+noise Average
Illegal Activitiy 97 16.49 - 51.55 80.41 49.48 15.46 - 73.20 84.54 57.73
Hate Speech 163 31.29 - 32.52 62.58 42.13 25.15 - 53.37 64.42 47.64
Malware Generation 44 75.00 - 31.82 70.45 59.09 77.27 - 52.27 59.09 62.87
Physical Harm 144 69.44 - 54.17 86.81 70.14 59.03 - 66.67 77.78 67.82
Economic Harm 122 22.95 - 22.95 36.07 27.32 28.69 - 28.69 35.25 30.87
Fraud 154 57.14 - 51.95 76.62 61.90 70.13 - 65.58 80.52 72.07
Sex 109 27.52 - 15.60 32.11 25.07 36.70 - 24.77 27.52 29.66
Privacy Violence 139 52.52 - 68.35 82.73 67.86 56.83 - 76.26 85.61 72.90
Average - 44.04 - 41.11 65.97 50.37 46.15 - 55.10 64.34 55.19